OpenStudy (anonymous):

23.6,-45.2,50,-0.5,0.3,3/5,-4/5 LEAST TO GREATEST?

OpenStudy (anonymous):

convert the fractions into decimals... so 3/5 = 0.6 -4/5 = -0.8 SO, that gives us: -45.2, -.8, -.5, 0.3 , 0.6 , 23.6, 50